The reaction SO2Cl2(g)----->SO2(g) + Cl2(g) is first order inSO2Cl2. Using the following kineticdata, determine the magnitude of the first-order rateconstant?
Time(s)                       Pressure SO2Cl2 (atm)
0                                          1.00
2,500                                   0.947
5,000                                   0.895
7,500                                   0.848
10,000                                 0.803
Answer #1
Graph between time and pressure can be plotted as follows,
                                
This is the graph for the first order reaction. Equationcorresponding to the graph is y = - 2.0 x 10-5 x-0.0001
Here slope is nothing but rate constant. ∴K =2.0 x 10-5 s-1
Answer #2
A = A0*exp(-kt)

ln(A) = ln(A0) - kt

Plot ln(A) vs t, slope of line is -k

plot

t ln(Pressure)
0 0
2500 -0.05
5000 -0.11
7500 -0.16
10000 -0.22

slope = -2.20e-5

Thus, k = 2.20 x 10^-5
Answer #3
            SO2Cl2(g)-----> SO2(g) +Cl2(g)
          Theintigrated form of rate law for the first order reactionis
          ln [A]t = -kt + ln[A]o
     
          This equation can also be written for the pressures.
           from the given data t = 2500 s
          Initial presure of SO2Cl2 is = 1.00 atm
           pressure after 2500 s is 0.947 atm
    
                ∴ ln 0.947 atm = - k * 2500 sec + ln1 atm
                ∴ rate constant k = 2.178 x 10-5s-1